Simplify ( sqrt of 7 + 4)(sqrt of 7 - 4)
Answer by mathchemprofessor(65)

Simplify ( sqrt of 7 + 4)(sqrt of 7 - 4)
=(7^1/2 +4)(7^1/2 -4)=(7^1/2)^2-4^2=7-4^2=7-16=-9